OpenStudy (anonymous):

help me plzz Reduce to lowest terms. (3m^2-3n^2)/(6m+6n)

OpenStudy (anonymous):

factor and cancel

OpenStudy (anonymous):

\[3m^2-3n^2=3(m^2-n^2)=3(m+n)(m-n)\] for the numerator

OpenStudy (anonymous):

um idk

OpenStudy (anonymous):

help me